Reports 1-1 of 1 Clear search Modify search
AdV-ISC (Commissioning up to first full interferometer lock)
boldrini, valentini - 0:05 Thursday 29 July 2021 (52691) Print this report
ISC shift: carm null robustness study

The goal of the shift was to study and improve the robustness of the ITF lock ad carm offset null, by studying the effects diff+ drift control, PR AA loop, UGF servos and eventually the TCS systems.

During our shift we realigned the green laser in the West arm. The North was also slightly misaligned, not enough to prevent the lock after increasing the reflection DC lock threshold to avoid having it unlocking every few minutes.

We proceeded towards carm null slowly, finely adjusting gains and phases along the way. In particular, we tuned the phase for B2_169MHz in every step of the way (in particular at 300mW and higher powers) in order to strike the best balance between reducing the coupling between PRCL/SRCL and MICH, by minimizing the magnitude of their corresponding UGF monitor lines on the latter error signal.

The first two times we reached carm null, we unlocked for a 70 Hz oscillation on DARM (Fig.1-2, row 2, plots 5-6) that we solved by reducing its gain by a 30% factor, and this problem did not manifest again for the rest of the shift.

We mainly investigated the behaviour of the quadrants used for diff+, B1p_QD2, at carm null. As also noted in the previous shift (52679) we noticed that at carm null the quadrants galvo loops start behaving erratically: the B1p QD2 galvos (Fig.3-4, row 4, plots 6-7) sometime open and the shutter closes as soon as there is any power spike in B1p. At the same time, the B2 QD2 quadrant used for PR AA galvo corrections perform a large jump (Fig.3-4, row 5, plots 1-2). Simultaneous jumps can be observed on the camera position of B2 (Fig.3-4, row 5, plots 4-5). Fig.7 shows these in detail, in particular that the jumps on the galvos loops happen exactly when one goes to carm null.

We decided to keep the diff+ closed up to 300mW, where we open it and let it be. Since the PR AA loop error signals degrade at carm null (see fig 4), possibly due the jumps observed on the B2 beam, we decided to keep it in drift control (Fig.5, row 4, plots 2-3) at carm null.

Furthermore, we noticed that DARM UGF drops significantly after reaching carm null, therefore we enabled the UGF servos for PRCL, MICH and DARM with target UGFs 45, 15, 40 respectively.

These interventions all togheter allowed for a lock in carm null marginally more stable than usual, lasting 5 minutes (last lock in Fig.8). For this last attempt, we also tuned the phase of B2_169MHz in carm null (Figs.9-10). As one can see in the picture, minimizing PRCL line on I significantly increases SRCL contribution. We chose to prioritize minimizing PRCL coupling with MICH.

We leave the arms locked on the IR.

Images attached to this report
Search Help